# Unveiling Challenges for LLMs in Enterprise Data Engineering

### v1

@@ 5.2. Experiments & Results (Case Study 1)
To empirically examine the challenges that arise when solving enterprise data engineering tasks with LLMs, we first conduct experiments for our first case study with the payment-to-invoice matching scenario, which allows us to show many of the issues emerging from challenges C5, C7, and C8 discussed before. Experiments for the second case study, which focuses on challenge C6 (compound tasks), follow in the next subsection. For our experiments, we are using a dataset based on real-world customer data. Moreover, we pre-processed the dataset so that we can systematically include the challenges discussed before and analyze the performance when incrementally adding these challenges (e.g., with and without 1:N matches, with and without data errors).
@@
For the analysis, we use a dataset of invoices and payments following the characteristics of the actual data from an enterprise software system and mirroring many of the challenges described in C5, C7, and C8. Our dataset contains 15,521 invoices and 12,332 payments, and we experiment on 790 matching pairs and 1,210 pairs that do not match. We work with the data in two formats: (1) one format where each entity is represented by multiple tables and (2) another format where it is represented as one view that combines all tables, which is closer to the academic setting. Moreover, we formulate the entity matching task as a binary classification between rows similar to existing literature (Narayan et al., 2022; Peeters and Bizer, 2023). As such, we prompt the LLM to decide if two table rows (one payment and one invoice) match. For this experiment, we use a few-shot approach (as zero-shot performance was really low for enterprise data) and include one positive and one negative example (a match and a non-match) in the prompt. More examples did not improve the accuracy significantly. Afterward, we make the task incrementally more complex.
@@
Exp. 4: Increasing task complexity. In the following, we explain the different scenarios with results provided in Table 5:
1. (clean) First, we run the models on clean 1:1 matches using the single table views, which is closest to the academic setting. In this setting, the payment’s memo line includes the correct reference numbers, the payment amount is exactly as stated in the invoice, and the customer name is also identical. As a result, the models reach very high F1 scores in this setup, which we can see in Table 5 (Clean). Next, we incrementally increase the difficulty of matching payments to invoices based on the challenges we observe in enterprise data.
2. (+ errors) As mentioned before in C8, data errors in enterprises are different from academic datasets, making the matching harder. To show this, we add representative errors and discrepancies from real data to the source and target data that typically occur in real-world customer data, such as missing or additional digits in the reference numbers or (minor) discrepancies in the paid amount. As shown in Table 5 (+ Errors), these seemingly small inconsistencies already cause a noticeable drop in accuracy, highlighting the sensitivity of the models to such errors. We further analyze this case below in Exp. 5, which showcases that LLMs have real problems understanding the semantics of enterprise data.
3. (+ multi-matches) Next, we make the matching even harder and focus on multi-match cases as explained in C7, where one payment pays multiple invoices or multiple payments together pay one invoice, which is a setting that is already closer to real-world scenarios found in enterprises. As shown in Table 5 (+ Multi-Matches), this vastly increases the task’s difficulty. To understand the root cause of this drop in F1 scores, we further investigate the precision and recall to understand the high deviation in F1 scores between the models (see Figure 6). We find that while all models achieve very high precision, differences in F1 scores are primarily driven by variations in recall. This indicates that all of the models take a rather cautious approach, only predicting matches when they are highly certain. As a result, they prioritize high precision in matching and therefore miss many correct matches. While this high precision is sometimes preferred over high recall but low precision, the high rate of missed matches causes high manual efforts for enterprises to find the missing matches.
4. (+ multiple tables) Finally, to show the effect that business entities often span multiple tables as mentioned in C5, we represent invoices using multiple tables (as in the original SAP schema) instead of the single flat table view. While metadata about each invoice document is stored in one table, specific information like the amount and due date are stored in a second table, and information about customers is stored in another separate customer table. In this scenario, which now represents the actual enterprise task in its entire complexity, all LLMs see large performance decreases compared to the previous experiments, with the lowest performance in GPT-4o-Mini, indicating that the models have difficulties working with the complex data structures often used in enterprises.
Exp. 5: Data errors amplify task complexity? Our second experiment analyzes the impact of different error types on complex enterprise tasks like payment-to-invoice matching, showcasing some interesting findings which relate the task to typical data challenges in enterprise scenarios. In this experiment, we insert individual classes of errors into the clean data of Exp. 4 to study their effects. We focus on four typical cases: (1) we deduct up to USD 0.1 from the amount paid, (2) we remove or change digits in the assignment or (3) in the billing number, and (4) we slightly vary the partner name, for example with abbreviations like KL Technologies instead of Kim & Lee Technologies. The results for the different error classes shown in Figure 7 indicate that even minor discrepancies lead to performance degradation of the LLMs. Interestingly, errors in numerical fields only cause a relatively modest performance drop (around 5%), while discrepancies in business partner names result in a much more severe decline. This suggests that for enterprise data, the LLMs rely heavily on the availability of non-erroneous textual fields for matching, highlighting their inability to understand numerical and symbolic data, which is quite prominent in enterprise data as discussed in Section 4.1.
5.3. Experiments & Results (Case Study 2)
To investigate the impact of the compound tasks stated in C6, we now conduct several experiments based on our second case study (database integration) introduced at the beginning of this section.
Setup.
Similar to before, we are using a dataset based on real-world data for data integration from two companies. Company A’s database consists of a single table with 15 columns, selected as a subset of the SAP general customer master data table KNA1 with columns such as KUNNR, NAME1, and ERDAT. Company B’s database consists of two tables with a foreign key relationship, where the tables have self-descriptive column names such as Organization Name, Address, and Tax Number. We experiment with datasets of different sizes, from 50 to 300 customers, to analyze the effects of scaling the data. The customers in the two databases have an overlap of 60% (i.e., 40% of the customers exist only in one of the databases). To evaluate the predicted result table, we compare it with the ground truth result table of integrated customers and compute the cell-level accuracy (i.e. the fraction of correct cells).
@@ 5.3. Experiments & Results (Case Study 2)
Exp. 6: Error propagation in compound tasks. In our first experiment for this case study, we aim to analyze the effect of error propagation in compound tasks with multiple steps. For this, we formulate the task of customer data integration as a sequence of the following three steps, visualized in Figure 2 (center):
@@
1. (schema matching) To integrate the data, we first perform schema matching to find the mapping between the columns of Company B’s database to the columns in Company A’s database. As in existing research, we formulate the task as a binary classification (Narayan et al., 2022; Parciak et al., 2024). Challenges arise because, for example, the address is represented in a single column in Company B, but is split into multiple fields in Company A. We measure the accuracy of this step as the fraction of correctly found matches for the columns in Company A since the output of the schema matching step is used to transform Company B’s database into Company A’s schema.
@@
2. (entity matching) To merge the data correctly, we perform entity matching to find out which customers have entries in both companies’ databases. Similar to the existing literature (Narayan et al., 2022; Peeters and Bizer, 2023), we formulate entity matching as a binary classification task: Given a pair of rows, one from Company A’s table and one from Company B’s transformed table, the LLM determines whether they refer to the same customer.
@@
3. (data integration) For the actual data integration, the final customer table is created by using the LLM to automatically merge the duplicate records found in both databases into a single row and transform Company B’s records into the format of Company A.
To evaluate the impact of error propagation when chaining the tasks, we run each step twice: once based on the output of the previous pipeline step, and once in a standalone setting; i.e., with the correct results (without errors) of the previous step as input. The results of this experiment are shown in Figure 8, where the results with error propagation (called Pipeline) are shown as colored bars, and the results of the Standalone setting (without error propagation) are shown as stacked, colorless bars.
@@
For entity matching and data integration (the second and third steps in the pipeline), we observe that executing these tasks as a pipeline leads to a drop in accuracy compared to running them as standalone tasks. However, the decrease depends on the model. While for GPT-4o-Mini (smallest model), we see a severe drop due to error propagation, Claude and Llama see a much smaller drop.
@@
Interestingly, the performance between the entity matching and data integration step in the pipeline does not decline and actually slightly improves for some models. This shows a surprising effect for LLMs: While the models make errors that propagate through the pipeline, they can, in some cases, correct earlier mistakes. For instance, some models fail to correctly associate the PSTLZ field with the postal code during schema matching. However, when presented with three sample rows from Company A during data integration, they are able to correctly populate the schema of Company A with the addresses stored in Company B’s data.
@@
Overall, these results suggest that compound tasks clearly lead to a drop in accuracy due to error propagation, but the significance depends on the LLM. In some cases, the effects are even negligible.
Exp. 7: End-to-end task formulation. A high overhead in enterprise tasks stems from the fact that complex tasks must first be manually decomposed into individual steps by humans. Recent LLMs, however, can do reasoning to solve multi-step problems on their own. As such, one might ask the question if the manual decomposition of the problem is even needed, or if recent LLMs that are able to reason can break down the problems on their own?
@@
To show evaluate this, we let the LLMs directly integrate the two customer databases with a single prompting step and compare the results to the pipelined task formulation from Exp. 6. We experiment with two prompting strategies: (1) w/o Steps provides a textual description with instructions on how to perform the integration but does not mention how to break down the task, whereas (2) with Steps explicitly mentions the necessary intermediate steps (schema matching, entity matching, and data integration) in the prompt to provide a hint towards how to execute them.
@@
The results are shown in Figure 3. In this experiment, we include OpenAI’s o1 as an additional LLM with reasoning capabilities. We observe that most models perform better in the pipelined task formulation. Interestingly, only GPT-4o-Mini achieves better results end-to-end when compared to the pipelined execution of the experiment before. When comparing end-to-end w/o Steps and with Steps, we see mixed effects. While some LLMs benefit from describing the steps in the prompts, others (e.g., o1 and Claude) do not.
Exp. 8: End-to-end scaling. While the end-to-end formulation of compound tasks is attractive as it does not require manual task decomposition, it necessitates that the full tables are handed in as input to the LLM because some steps (e.g., a matching step between tables) might need access to all tuples.
@@
As a last experiment in our customer integration case study, we investigate the effects of scaling the number of customers in the company databases. However, in the end-to-end task formulation, this leads to longer prompts that quickly exceed the available context windows. Among the models evaluated in this work, only OpenAI’s o1 reasoning model provides a sufficiently high output token limit (100K tokens), which is necessary to integrate more than 100 customers in an end-to-end manner.
@@
Figure 10 presents the results of scaling the total number of customers from 50 up to 300. First, we observe a drastic decline in performance as the number of customers increases, suggesting that the model struggles with long-range dependencies and finding matches in large datasets. Moreover, it often fails to merge duplicate customers. When using 300 customers, the model even responds with ”Sorry, but I can’t fulfill that”. Overall, the ability of LLMs to handle enterprise-scale tasks end-to-end remains limited. In contrast, a formulation that breaks up compound tasks into individual steps, as in our previous experiment, can, in principle, scale to larger tables as tuples are handed in one by one (e.g., comparing all customers of databases A and B individually for matching).
@@ 5.4. Discussion
In our experiments, we experienced that automating enterprise tasks end-to-end with LLMs is extremely hard: First, we still need to rely on human effort as end-to-end approaches with LLMs have been found to be very brittle and not scalable. This is consistent with Ashury-Tahan et al. (2025), who found that even strong models often fail to perform robustly on complex tabular data tasks. In addition, we encountered problems in the output of some task formulations, like end-to-end, where the output is a table with multiple rows, since models often generate varying numbers of columns for individual rows. This typically does not happen with task formulations where the output is just a single row.
Second, much of the complexity of enterprise tasks is inherent to business needs and cannot be avoided, such as the need to deal with multi-match cases in entity-matching. Furthermore, we found that reducing complexity by working with database views improved the results. However, constructing such a view causes additional manual overheads as business objects in enterprises can be composed of tens to even hundreds of tables. Finally, some challenges like error propagation are interestingly not as severe as initially thought. Overall, however, given the current abilities, LLMs are still not sufficient to achieve the level of performance required for enterprise-scale data engineering.
@@
To better handle the complexity of enterprise tasks, we think that LLMs on their own will not be sufficient since much of the complexity cannot be simplified, as discussed earlier. Instead, complex tasks need to be approached with systems that combine LLMs with other methods while also designing effective human-LLM interactions. Wornow et al. (2024) propose demonstrating workflows to foundation models, which is promising for automating frequently occurring tasks. However, such solutions still require manual overheads. Another promising direction is leveraging LLMs to first create structured plans for approaching compound tasks similar to recent advancements in using LLMs for query planning (Urban and Binnig, 2024; Liu et al., 2025) and then execute these plans step-wise using existing approaches (e.g., heuristics for schema matching) or even the LLM itself.

### v2

@@ 5.2. Experiments & Results (Case Study 1)
To demonstrate how these challenges affect data engineering with LLMs, we first conduct experiments for our first case study (entity matching). Experiments for our second case study (database integration) follow in the next subsection.
@@
For our first case study, we use a dataset of payments and invoices following the characteristics of the actual data from an enterprise system and reflecting the challenges described above (see Appendix A.2 in our extended technical report). The dataset contains 15,521 invoices and 12,332 payments. Following existing literature (Narayan et al., 2022; Peeters and Bizer, 2023), we formulate the entity matching task as a binary classification where the LLM decides if two entities (one payment and one invoice) match; i.e., if the payment pays for the invoice. Note that some of these combinations are part of multi-match cases where one payment pays for multiple invoices or vice versa. To keep the costs tractable, we use a subset of the dataset consisting of 790 payment-invoice combinations where the payment does pay for the invoice and 1,210 combinations where it does not.
@@
For each combination, we prompt the LLM to predict if it is a match or a non-match. Furthermore, we experiment with presenting the invoices in two formats to the LLM: one format where each invoice is represented by multiple separate tables, and another format where each invoice is represented as one row in a flat view that combines all relevant columns from these tables, mimicking the traditional entity matching setting. We use a few-shot approach that includes one positive and one negative example in the prompt, as zero-shot performance was low on enterprise data. More examples did not improve the accuracy much. Starting from this simple setup, we make the task incrementally more complex.
Exp. 5: Increasing task complexity. In the following, we explain the different scenarios with results provided in Table 6:
1. (Clean) First, we evaluate on clean 1:1 matches using the flat table views described above, which is closest to the academic setting. In this setting, the payment’s memo line includes the correct reference numbers, the payment amount is exactly as stated in the invoice, and the customer name is also identical. As shown in Table 6 (Clean), all models reach very high F1 scores in this setup. Next, we incrementally increase the difficulty based on the challenges we observe in the enterprise scenario.
2. (+ Errors) As described in C8, data errors in enterprises can differ from those in academic benchmarks, making the matching harder. To demonstrate this, we add representative errors and discrepancies that typically occur in real-world customer data to the payments and invoices, such as missing or additional digits in the reference numbers or minor discrepancies in the paid amount. As shown in Table 6 (+ Errors), these seemingly small inconsistencies already cause a noticeable drop in accuracy, highlighting the sensitivity of the models to such errors. We further study this effect in Exp. 6, showcasing that LLMs have substantial problems understanding the semantics of enterprise data.
3. (+ Multi-Matches) Next, we make the matching even harder by focusing on the multi-match cases described in C7, where one payment pays for multiple invoices or multiple payments together pay for one invoice, a setting that is already closer to the real-world scenarios found in enterprises. As shown in Table 6 (+ Multi-Matches), this vastly increases the task’s difficulty, since amounts are split between multiple payments and payments might include multiple reference numbers for different invoices. To understand the root cause of the performance decrease and the high deviation between the models, we further investigate the precision and recall (see Appendix A.2 in our extended technical report). We find that while all models achieve a very high precision close to , the differences in F1 scores are primarily driven by variations in recall. This indicates that all models take a rather cautious approach, predicting matches only when they are highly certain. As a result, they prioritize high precision in matching and therefore miss many correct matches. While this high precision is sometimes preferred over high recall but low precision, the high rate of missed matches causes substantial manual efforts for enterprises to find the missing matches.
4. (+ Multiple Tables) Lastly, to show the impact of business entities spanning across multiple tables as described in C5, we represent invoices using multiple separate tables (as in the original SAP schema) instead of a single flat table view. While metadata about each invoice is stored in one table, specific information like the amount and due date is stored in a second table, and information about customers is stored in another separate customer table. In this scenario, which now closely resembles the actual enterprise task in its full complexity, all models see large performance decreases compared to the previous experiments, indicating that the models have difficulties working with the complex data structures used in enterprises.
Exp. 6: Data errors and task complexity. Our second experiment analyzes the impact of different error types on enterprise tasks like payment-to-invoice matching, highlighting typical data challenges that appear in enterprise scenarios. We insert individual types of data errors into the clean data from Exp. 5 to study their effects in isolation. We focus on four typical cases: (1) we deduct up to USD 0.1 from the paid amount, (2) we remove or change digits in the assignment or (3) in the billing number, and (4) we slightly vary the partner name, for example with abbreviations like KL Technologies instead of Kim & Lee Technologies. The results shown in Figure 6 indicate that even minor discrepancies can lead to performance degradations. Interestingly, errors in the numerical fields cause only modest performance declines, whereas discrepancies in the business partner names result in much more severe performance drops. This suggests that LLMs rely heavily on the availability of non-erroneous textual fields for entity matching, highlighting their inability to understand the complex numerical and symbolic values prominent in enterprise data.
5.3. Experiments & Results (Case Study 2)
To study the compound tasks described in C6, we now conduct experiments based on our second case study (database integration).
Setup.
We use a dataset for the integration of two companies’ customer databases that is based on real-world data (see Appendix A.3 in our extended technical report). Company A’s database consists of a single table with 15 columns, selected as a subset of the SAP customer master data table KNA1 with columns such as KUNNR, NAME1, and ERDAT. Company B’s database consists of two tables connected by a foreign key relationship, with descriptive column names such as Organization Name, Address, and Tax Number. We experiment with datasets of different sizes, from 50 to 300 customers, to analyze the effects of the data scale. The customers in the two databases have an overlap of 60%, so 40% of the customers exist only in one of the databases. To evaluate the predicted result table, we compare it with the ground truth result table of integrated customers and compute the cell-level accuracy (i.e. the fraction of correct cells).
@@ 5.3. Experiments & Results (Case Study 2)
Exp. 7: Error propagation in compound tasks. In our first experiment, we analyze the effect of error propagation in compound tasks with multiple steps. We formulate the database integration task as a sequence of three steps, as visualized in Figure 2 (center):
@@
1. (Schema Matching) To integrate the data, we first perform schema matching to map the columns of Company B’s schema to the columns of Company A’s schema. Following existing research, we formulate the task as a binary classification between columns (Narayan et al., 2022; Parciak et al., 2024). Challenges arise because, for example, the address is represented by a single column at Company B, but split into multiple columns at Company A. We measure the accuracy of this step as the fraction of correctly found matches for the columns in Company A, since the output of the schema matching step is used to transform Company B’s database into Company A’s schema.
@@
2. (Entity Matching) Next, we perform entity matching to determine which customers have entries in both companies’ databases. Similar to existing literature (Narayan et al., 2022; Peeters and Bizer, 2023), we formulate entity matching as a binary classification task: Given a pair of rows, one from Company A’s table and one from Company B’s transformed table, the LLM determines whether they refer to the same customer.
@@
3. (Data Integration) To produce the integrated customer table, we use the LLM to merge the duplicate customer records found in both databases into a single row and to transform Company B’s remaining customer records into the format of Company A.
@@
To measure the impact of error propagation when chaining the tasks, we run each step twice: once in a pipeline using the output of the previously executed step as input, and once in a standalone setting using the ground-truth output of the previous step as input. Figure 7 shows the results of this experiment. For the second (entity matching) and third (data integration) step of the pipeline, we observe that executing them in a pipeline leads to a decrease in accuracy compared to running them as standalone tasks. However, the decrease depends on the model, with GPT-4o-Mini seeing a more severe decrease than Claude 3.5 Sonnet and Llama 3.1 Instruct.
@@
Interestingly, the accuracy between the entity matching and data integration steps in the pipeline does not decline and even slightly improves for some models. This shows a surprising effect for LLMs: While the models can make mistakes that propagate through the pipeline, they can, in some cases, also correct earlier mistakes. For example, some models fail to correctly associate the LAND1 field with the address during schema matching. However, when presented with three sample rows from Company A during data integration, they correctly populate the LAND1 field with the country mentioned in the address stored in Company B’s data.
@@
Overall, these results suggest that error propagation can lead to a decrease in accuracy for compound tasks, but the significance depends on the LLM. In some cases, the effects are negligible.
@@
Exp. 8: End-to-end task execution. A substantial overhead in enterprise scenarios stems from the fact that complex tasks must be manually decomposed into smaller steps. Recent LLMs, however, can use reasoning to solve multi-step problems on their own. One might now ask the question if the manual decomposition of the task is even needed, or if reasoning LLMs can break down the task on their own. To answer this question, we use the LLMs to directly integrate the two customer databases in a single LLM call and compare the results of this end-to-end execution to the pipelined execution from Exp. 7. We experiment with two prompting strategies: w/o steps provides a textual description with instructions on how to perform the integration but does not mention how to break down the task, whereas with steps explicitly mentions the necessary steps (schema matching, entity matching, and data integration) in the prompt to provide a hint towards how to execute them.
@@
Figure 4 shows the results of this experiment, including o1 from OpenAI as an additional LLM with reasoning capabilities. Most models perform better in the pipelined task execution, with only GPT-4o-Mini performing better in the end-to-end execution. When comparing the end-to-end execution w/o steps and with steps, we see mixed results. Whereas some LLMs benefit from describing the steps in the prompts, others do not.
@@
Exp. 9: End-to-end scaling. While the end-to-end execution of compound tasks is attractive because it does not require manual task decomposition, it requires the full tables to be provided as input to the LLM because some steps like entity matching might need access to all tuples. This poses a problem, as larger tables lead to longer prompts that quickly exceed the LLM’s context window. As a last experiment in our database integration case study, we study the effects of scaling the number of customers in the company databases. Among the models evaluated in this work, only OpenAI’s o1 provides a sufficiently large output token limit (100K tokens) to integrate more than 100 customers in an end-to-end manner.
@@
Figure 9 shows the results of scaling the total number of customers from 50 up to 300. We observe a drastic decline in performance as the number of customers increases, suggesting that the model struggles with long-range dependencies and finding matches in large datasets. When using 300 customers, the model even responds with ”Sorry, but I can’t fulfill that”. Overall, LLMs remain limited in their ability to execute tasks at enterprise-scale in an end-to-end manner. In contrast, breaking up compound tasks into individual steps can, in principle, scale to larger tables, as tuples are processed one by one (e.g., by comparing all customers of databases A and B individually for entity matching).
@@ 5.4. Discussion
Through our two case studies, we have demonstrated that automating enterprise tasks with LLMs remains extremely hard. First, we still need to rely on human effort since end-to-end execution with LLMs is brittle and does not scale well. This is consistent with Ashury-Tahan et al. (2025), who found that even strong models often fail to perform robustly on complex table tasks. We also encountered problems in the output of some task formulations. For example, during end-to-end execution, where the output is a table with multiple rows, LLMs often generate varying numbers of columns for different rows.
Second, much of the complexity of enterprise tasks is inherent to business needs and cannot be avoided, such as having to deal with multi-match cases in entity matching. While reducing complexity by working with database views did improve results, constructing such views causes additional manual overheads since business objects in enterprises can be composed of tens or even hundreds of tables. Finally, some challenges like error propagation are not as severe as initially thought. Nevertheless, given their current abilities, LLMs are still not sufficient to achieve the level of performance required for enterprise-scale data engineering.
@@
We believe that LLMs out-of-the-box will not be sufficient to address the complexity of enterprise tasks because much of it is inherent to enterprise needs. Instead, complex tasks call for carefully-designed systems that use LLMs as reasoning or tool-calling agents while effectively incorporating human-LLM interactions. Wornow et al. (2024) propose demonstrating workflows to foundation models, which is promising for automating frequently occurring tasks but still requires manual overheads. Another promising approach uses LLMs to create structured plans for compound tasks similar to recent advancements in LLMs for query planning (Urban and Binnig, 2024; Liu et al., 2025b) before executing these plans step-by-step using existing approaches (e.g., heuristics for schema matching) or even the LLM itself. This direction is reinforced by our findings: our experiments show that the stepwise execution of compound tasks improves performance over end-to-end execution while also enabling the LLM to correct earlier mistakes, for example by using information from subsequent steps.